OpenStudy (anonymous):

Find the intersection points of the graph of y=x^2 and 2+y=-3x

OpenStudy (anonymous):

solve for y: y=-3x-2 then equate: -3x-2=x^2 x^2+3x+2=0 (x+2)(x+1)=0 x=-2 or x=-1 then you can back sub for the y-values

OpenStudy (anonymous):

y=-3(-2)-2=6-2=4 --> (-2,4) y=-3(-1)-2=3-2=1 --> (-1, 1)

OpenStudy (anonymous):

or y=(-2)^2=4 --> (-2,4) y=(-1)^2=1 --> (-1,1)